# Two failure modes of deep transformers and how to avoid them: a unified theory of signal propagation at initialisation

### v2

@@ 1 Introduction
Finding the right initialisation for a neural network is a key challenge facing every practitioner before training. Choosing the right scale for the initial weights is critical for maintaining information flow during the forward and backward passes, which in turn ensures smooth training and good final performance. In fully connected networks, a series of works established a clear picture of signal propagation and trainability at initialisation by computing how the similarity of two inputs evolves as they propagate through the same network at initialisation [1, 2, 3, 4, 5, 6], ultimately establishing the optimal initialisation at the “edge of chaos” [1].
@@
In Transformers [7], where fully-connected layers alternate with self-attention layers [8], the key quantity for measuring information flow through the network is the similarity between tokens in a sequence as it propagates through the network. Signal propagation faces additional challenges due to two key failure modes of self-attention layers. The first is rank collapse, where self-attention maps all input tokens to identical representations, producing an output matrix of rank one. This phenomenon manifests as the attention pattern shown in Fig.˜1(a). [9] demonstrated that networks composed solely of self-attention layers will inevitably collapse any input sequence into uniform token representations at a rate that is double exponential in the number of layers. This rank collapse fundamentally destroys input sequence information and impedes effective training by inducing vanishing gradients [10]. Entropy collapse represents the second failure mode, where queries attend to only a small, frozen subset of tokens irrespective of the input, leading to low Shannon entropy of the attention distribution (hence the name) and, more importantly, unstable training [11]. An attention matrix exhibiting this pathology is shown in Fig.˜1(b).
@@
Figure 1: Two failure modes of Transformers at initialisation, and how to avoid them. (a) Rank collapse occurs when the self-attention layer attends uniformly to all tokens, mapping all input tokens into the same output token. (b) Entropy collapse is a regime of highly saturated attention matrices which attend to random, semantically meaningless patterns, leading to training instability [11]. (c) Trainability diagram for a 60-layer BERT transformer, obtained from our analytical theory of signal propagation, see Section˜5. Depending on the strength of the self-attention residual connections $\alpha_{\mathrm{SA}}$ (Eq.˜5) and the scale of initial key and query weights $\beta$ (Eq.˜8), we delineate the three regimes of rank collapse, entropy collapse, and the regime where the Transformer is trainable (blue). (d) Average cosine similarity between token embeddings of a sequence taken from the TinyStories dataset as it propagates through the layers of a vanilla BERT model for different self-attention residual strengths; empirical measurements (dots) closely follow theoretical predictions (solid lines). Sufficiently large residual connections $\alpha_{\text{SA}}$ are key to preventing the similarity between tokens from becoming unity, which would indicate rank collapse. (e) Test loss of a 60-layer BERT model on TinyStories for two initialisations from each regime. Models suffering from rank or entropy collapse at initialisation fail to train, as predicted by theory. Full experimental details in Section˜A.1.
@@
Previous work has highlighted the importance of skip connections in mitigating rank collapse [9, 10, 12], and [11] suggested a modification of the self-attention layer that helps avoiding entropy collapse. However, a quantitative, unified description of how these two distinct phenomena emerge has been lacking. Our Result˜1 fills this gap, providing clear guidance on how to jointly avoid both issues by appropriately choosing the strength of residual connections and the scale of initial weights.
@@
The main challenge in studying information propagation in Transformers arises from the self-attention layers. Previous works either assume uniform attention [10] or approximate the average behaviour of a self-attention layer by taking expectations separately over the numerator and denominator of the softmax [13] – a strong simplification that fails the behaviour of large initial query and key weights, which is responsible for entropy collapse.
In contrast, we analyse the standard softmax transformer in the complementary limit of infinitely long sequences by leveraging tools and concepts from statistical physics. Our theoretical framework provides a unified explanation for – and a practical solution to – the emergence of both failure modes observed in practice: rank collapse and entropy collapse.
Our main contribution
@@
is an analytical theory of signal propagation in deep transformers with self-attention layers, skip connections, layer normalisation, and MLPs. Our theoretical framework yields a simple algorithm to compute the evolution of the typical overlap between token embeddings as sequences propagate through deep, off-the-shelf transformers at initialisation. This enables us to construct trainability diagrams such as the one shown in Fig.˜1(c) for a 60-layer BERT-style transformer. By varying the strength the scale of initial key and query weights $\beta$, Eq.˜8, and the strength of the skip connections of the self-attention layer $\alpha_{\text{SA}}$, Eq.˜5, we identify three regimes for signal propagation: entropy collapse dominance (red), rank collapse dominance (yellow), and a trainable regime characterised by small initial weights and strong skip connections (blue). The critical threshold for query/key variance, $\beta_{c}$, emerges as a global property, while the critical threshold $\alpha_{c}$ for residual strength depends on network depth: for any given model depth, our theory predicts the minimum residual strength required to guarantee signal propagation and ensure trainability.
We validate our theory in two ways.
@@
In Fig.˜1(d), we show that our theory (solid lines) accurately predicts the average cosine similarity between tokens (averaged over all token pairs in a sequence) when propagating sequences from the TinyStories dataset [14] through a vanilla BERT model [15] at initialisation (dots). In Fig.˜1(e), we show the test loss of a 60-layer BERT model on TinyStories for two initialisations from each regime; models suffering rank or entropy collapse at initialisation indeed fail to train, as the trainability diagram predicts. We provide further applicaitons of our theory in Section˜5.
@@
  • We identify a sharp phase transition in the behaviour of self-attention in the limit of infinite sequence length, using tools from statistical physics [16]. Below a critical value $\beta_{c}$ of the initial scale $\beta$ of query/key weights, Eq.˜8, attention matrices are approximately uniform, while for $\beta>\beta_{c}$, they are highly highly saturated, causing rank and entropy collapse, respectively, see Result˜1.
  • Our results are exact in the limit of very large sequences. We calculate finite-size corrections that are essential for obtaining precise, practical predictions, see Section˜3.2.2.
@@
  • We integrate our results on self-attention and normalisation layers with existing results on MLPs to derive simple algorithms like Section˜5 that yield trainability diagrams like the one shown in Fig.˜1d, see Section˜4.
@@
Further related works.
[17, 18] developed an elegant mathematical framework that models transformers as interacting particle systems, revealing emergent clustering phenomena in token representations.[19] determines Bayes-optimal generalisation error for attention-indexed models at finite sequence length $T$, while we consider the complementary limit of $T\to\infty$. [20] analyses mean-field training dynamics in various limits: infinite query/key dimensions, infinite heads, and infinite depth, leveraging dynamical mean field theory to derive feature and gradient kernels. [21] provide a description of the limiting distribution of the covariance matrix of token embeddings in the proportional limit of infinite width and depth for a modified Transformer architecture; we revisit their result in Section˜5. We explore the complementary limit of infinite sequence length, providing a framework to investigate signal propagation in standard Transformers as well as in a range of architectural variations.
@@ 2 Setup and notation
Goal
We consider a vanilla Transformer encoder [7] that processes sequences $\{X_{t}\}_{t=1,\ldots,T}$ of $T$ tokens embedded in a $d$-dimensional space. We analyse signal propagation through a complete Transformer block comprising self-attention, residual connections, layer normalisation, and a feed-forward MLP.
Layer Norm.
We consider layer normalisation [22] which centers each token embedding and rescales it by its standard deviation; for simplicity, we omit the affine transformation:
@@
Self-Attention.
Given the query, key and value weight matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$, the attention score between tokens $X_{t}$ and $X_{t^{\prime}}$ is
@@
MLP.
Each embedding $X_{t}$ is processed independently by a shared two-layer feed-forward network with hidden dimension $d$ and non-linearity $\phi$:
@@
Residual Connections.
We write the residual connections for the self-attention and MLP blocks as
@@
where $\alpha>0$ are the strenghts of the skip connections.
Notation
@@ 3 A theory for signal propagation in Transformers
Outline
The main challenge in analysing information flow through a transformer is handling the self-attention mechanism with its strong non-linearity and its normalisation step. In this section, we derive a theory for signal propagation by leveraging a formal similarity between self-attention and the Random energy model of [16], see Section˜3.1. Using tools and concepts from statistical physics, we give an asymptotically precise characterisation of signal propagation in self-attention layers in the forward pass in Section˜3.2. We analyse the backward pass in Section˜3.3. To complete the picture, we finally add the remaining elements of the transformer architecture in Section˜4, which yields Section˜5 to compute the trainability diagram of transformers.
@@ 3.1 Self-attention and the random energy model
The starting point of our analysis is an exact mapping of attention at initialisation to a model of disordered systems from statistical physics, called the Random Energy Model [16]. This mapping was previously identified by [23] in the context of associative memories; our novelty lies in leveraging it to provide an exact description of signal propagation.
The Random Energy Model (REM).
The REM is a simple model of a disordered system with $N$ spins $\mathbf{s}$ that has $\mathcal{O}(\exp(N))$ possible configurations. Each configuration has an energy $E(\mathbf{s})$ drawn at random from a Gaussian distribution with variance $\mathcal{O}(N)$ – hence the name Random Energy Model. The probability of each state is given by the Boltzmann distribution with inverse temperature $\beta>0$:
@@
Mapping of self-attention to the REM.
We initialise the query/key projection parameters i.i.d. with variance $\sigma_{Q}^{2}=\sigma_{K}^{2}=\nicefrac{{\sigma_{a}}}{{d}}$ so that the attention scores, defined in Eq.˜2, are normally distributed with mean 0 and variance $\sigma_{a}^{2}$. While both REM energies and self-attention scores $a_{tt^{\prime}}$ are Gaussian, the key difference between them is that attention scores are correlated. Averaging over the realisation of $W_{Q},W_{K}$, we get the following correlation structure (see Section˜B.1):
$\text{Cov}(a_{ts},a_{\tau\sigma})=\sigma_{Q}^{2}\sigma_{K}^{2}\,q_{t\tau}\,q_{s\sigma}.$ (7)
@@
Scaling of Query/Key weights.
In MLPs, the natural scaling for weight variance is $\mathcal{O}(1/d)$ [24], ensuring that neuron pre-activations remain $\mathcal{O}(1)$. For self-attention, the REM analogy suggests that the natural scaling for query and key weight variance should make the attention score variance $\mathcal{O}(\log T)$, where $T$ is the sequence length, to match the energy of a REM with $N$ degrees of freedom, whose energy typically scales as $\mathcal{O}(N)$ with fluctuations of $\mathcal{O}(\sqrt{N})$, while the partition sum runs over $\mathcal{O}(e^{N})$ configurations. Self-attention only sums over $T$ tokens in the sequence, suggesting that attention scores should scale like $\mathcal{O}(\sqrt{\log T})$. This motivates us to control the variance of attention scores via a constant inverse "temperature" parameter $\beta\in\mathbb{R}$ as:
@@
This scaling is, to our knowledge, novel and complements the existing literature on various reparametrisations of the infinite-width limit [25] by accounting for the infinite sequence length limit.
@@ 3.2 Forward Signal Propagation through a self-attention layer
3.2.1 Quantifying information propagation through a self-attention layer
Assuming i.i.d. token embeddings, standard concentration results in high dimensions imply that their norms concentrate, while pairwise overlaps scale as $O(d^{-1/2})$. Under self-attention, these tokens mix and their overlaps evolve as the sequence propagates in depth. Two key quantities to describe the evolution of token similarity are the overlap and cosine similarity matrices
which measure the degree of alignment between two tokens in a sequence. The third important quantity is the inverse participation ratio (IPR) of an attention row, which is defined for all $t\in T$ and $r\in\mathbb{N}$ as
@@ 3.2.1 Quantifying information propagation through a self-attention layer
For $r=2$, the IPR estimates the effective number of tokens receiving attention. A small IPR, scaling as $o_{T}(1)$, indicates attention is evenly spread out over keys, while an IPR that is non-vanishing indicates that only $O_{T}(1)$ keys are relevant, resulting in sparse self-attention. The connection with Shannon entropy is direct: the IPR captures the collapse of entropy to zero. Higher values of $r$ make the distinction between localised and delocalised attention vectors even sharper.
3.2.2 Main result
By tracking the average norm $\mathbb{E}\langle q_{tt}\rangle:=q$ and the average overlap $\mathbb{E}\langle q_{ts}\rangle:=p$, we describe the evolution of the average cosine similarity $\rho$ and characterize the average IPR of an attention row, which serves as an indicator of entropy collapse. We state here the result and provide a derivation in Section˜B.2.
@@ 3.2.2 Main result
Let $W_{Q}$ and $W_{K}$ be initialised with i.i.d. entries with variance $\sigma^{2}_{Q}=\sigma^{2}_{K}=\beta\sqrt{\log T}/d$, and $W_{V}$ with variance $\sigma_{V}^{2}=\sigma_{v}^{2}/d$. For a sequence with average token norm $q$ and average pairwise overlap $p$, define the critical initialisation scale $\beta_{c}(q,p)\equiv\sqrt{\frac{2}{q(q-p)}}$. In the limit of infinite sequence length $T\to\infty$, we then have that:
@@
Figure 2: Phase diagram for a single layer of self-attention. We use Result˜1 to plot the average cosine similarity between pairs of tokens after one layer of self-attention as a function of the query/key variance parameter $\beta$ and the input average cosine similarity $\rho$. (Left): Theoretical phase diagram obtained from Result˜1 (for $q=1,\,p=\rho$). For $\beta<\beta_{c}$, we observe a rank-collapse phase, where all input tokens map to a single output token and cosine similarity saturates at 1. For $\beta>\beta_{c}$, token diversity is preserved. (Right): Simulations with embedding dimension $d=512$ and sequence length $T=1024$ qualitatively reproduce the theoretical transition, with deviations attributed to finite-size effects, as discussed in Section˜3.2.2.
Definition of $\beta_{c}$ and finite-size effects
Key to our analysis is the convergence behaviour of the softmax normalisation, $Z(\beta)=\sum_{\tau=1}^{T}e^{a_{t}\tau}$, which depends on $\beta$ through $a_{t\tau}$, see Eq.˜2. We identify $\beta_{c}$ as the $\beta$ at which the law of large numbers for $Z$ breaks down, i.e. the $\beta$ for which $Z(\beta)$ does not concentrate around its mean as $T\to\infty$. However, as [26] pointed out, the central limit theorem for $Z$, which in addition stipulates that $Z$ converges to its mean with Gaussian fluctuations around $\mathbb{E}Z$ are Gaussian, already breaks down at $\tilde{\beta}_{c}=\beta_{c}/2$. In practice, $T$ is large but finite, so we expect our asymptotic predictions to hold for $\beta<\tilde{\beta}_{c}$, with subleading Gaussian fluctuations. Above $\tilde{\beta}_{c}$, these finite-size fluctuations become non-Gaussian, suggesting a crossover to the broken-LLN regime rather than a sharp phase transition exactly at $\beta_{c}$. This reasoning explains the discrepancy between theory and simulations observed in Fig.˜2 around the critical inverse temperature.

### v3

@@ 1 Introduction
Finding the right initialisation for a neural network is a key challenge facing every practitioner before training. Choosing the right scale for the initial weights is critical for maintaining information flow during the forward and backward passes, which in turn ensures smooth training and good final performance. In fully connected networks, a series of works established a clear picture of signal propagation and trainability at initialisation by computing how the similarity of two inputs evolves as they propagate through the same network at initialisation (Poole et al., 2016; Schoenholz et al., 2016; Xiao et al., 2018; Hanin and Rolnick, 2018; Hanin, 2018; Doshi et al., 2023), ultimately establishing the optimal initialisation at the “edge of chaos” (Poole et al., 2016).
@@
In Transformers (Vaswani et al., 2017), where fully-connected layers alternate with self-attention layers (Bahdanau et al., 2015), the key quantity for measuring information flow through the network is the similarity between tokens in a sequence as it propagates through the network. Signal propagation faces additional challenges due to two key failure modes of self-attention layers. The first is rank collapse, where self-attention maps all input tokens to identical representations, producing an output matrix of rank one. This phenomenon manifests as the attention pattern shown in fig.˜1(a). Dong et al. (2021) demonstrated that networks composed solely of self-attention layers will inevitably collapse any input sequence into uniform token representations at a rate that is double exponential in the number of layers. This rank collapse fundamentally destroys input sequence information and impedes effective training by inducing vanishing gradients  (Noci et al., 2022). Entropy collapse represents the second failure mode, where queries attend to only a small, frozen subset of tokens irrespective of the input, leading to low Shannon entropy of the attention distribution (hence the name) and, more importantly, unstable training (Zhai et al., 2023). An attention matrix exhibiting this pathology is shown in fig.˜1(b).
@@
Figure 1: Two failure modes of Transformers at initialisation, and how to avoid them. (a) Rank collapse occurs when the self-attention layer attends uniformly to all tokens, mapping all input tokens into the same output token. (b) Entropy collapse is a regime of highly saturated attention matrices which attend to random, semantically meaningless patterns, leading to training instability (Zhai et al., 2023). (c) Trainability diagram for a 60-layer BERT Transformer, obtained from our analytical theory of signal propagation, see section˜4. Depending on the strength of the self-attention residual connections $\alpha_{\mathrm{SA}}$ (eq.˜5) and the scale of initial key and query weights $\beta$ (eq.˜9), we delineate the three regimes of rank collapse, entropy collapse, and the regime where the Transformer is trainable (blue). (d) Average cosine similarity between token embeddings of a sequence taken from the TinyStories dataset as it propagates through the layers of a vanilla BERT model for different self-attention residual strengths; empirical measurements (dots) closely follow theoretical predictions (solid lines). Sufficiently large residual connections $\alpha_{\text{SA}}$ are key to preventing the similarity between tokens from becoming unity, which would indicate rank collapse. (e) Test loss of a 60-layer BERT model on TinyStories for two initialisations from each regime. Models suffering from rank or entropy collapse at initialisation fail to train, as predicted by theory. Full experimental details in section˜A.1.
1.1 Related Work
@@
Previous work has highlighted the importance of skip connections in mitigating rank collapse (Dong et al., 2021; Noci et al., 2022; Wang et al., 2024), and Zhai et al. (2023) suggested a modification of the self-attention layer that helps avoiding entropy collapse. However, a quantitative, unified description of how these two distinct phenomena emerge has been lacking. Our result˜1 fills this gap, providing clear guidance on how to jointly avoid both issues by appropriately choosing the strength of residual connections and the scale of initial weights.
@@
The main challenge in studying information propagation in Transformers arises from the self-attention layers. Previous works either assume uniform attention (Noci et al., 2022) or approximate the average behaviour of a self-attention layer by taking expectations separately over the numerator and denominator of the softmax (Cowsik et al., 2024) – a strong simplification that fails the behaviour of large initial query and key weights, which is responsible for entropy collapse. In contrast, we analyse the standard softmax Transformer in the complementary limit of infinitely long sequences by leveraging tools and concepts from statistical physics. Our theoretical framework provides a unified explanation for – and a practical solution to – the emergence of both failure modes observed in practice: rank collapse and entropy collapse.
Further related works. Geshkovski et al. (2023a; b) developed an elegant mathematical framework that models Transformers as interacting particle systems, revealing emergent clustering phenomena in token representations. Boncoraglio et al. (2025) determines Bayes-optimal generalisation error for attention-indexed models at finite sequence length $T$, while we consider the complementary limit of $T\to\infty$. Bordelon et al. (2024) analyses mean-field training dynamics in various limits: infinite query/key dimensions, infinite heads, and infinite depth, leveraging dynamical mean field theory to derive feature and gradient kernels. Noci et al. (2023) provide a description of the limiting distribution of the covariance matrix of token embeddings in the proportional limit of infinite width and depth for a modified Transformer architecture while we explore the complementary limit of infinite sequence length. We revisit their result in section˜5.
1.2 Contributions
@@
Our main contribution is an analytical theory of signal propagation in deep Transformers with self-attention layers, skip connections, layer normalisation, and MLPs. Our theoretical framework yields a simple algorithm to compute the evolution of the typical overlap between token embeddings as sequences propagate through deep, off-the-shelf Transformers at initialisation. This enables us to construct trainability diagrams such as the one shown in fig.˜1(c) for a 60-layer BERT-style Transformer. By varying the scale of initial key and query weights $\beta$, eq.˜9, and the strength of the skip connections of the self-attention layer $\alpha_{\text{SA}}$, eq.˜5, we identify three regimes for signal propagation: entropy collapse dominance (red), rank collapse dominance (yellow), and a trainable regime characterised by small initial weights and strong skip connections (blue). The critical threshold for query/key variance, $\beta_{c}$, emerges as a global property, while the critical threshold $\alpha_{c}$ for residual strength depends on network depth: for any given model depth, our theory predicts the minimum residual strength required to guarantee signal propagation and ensure trainability.
@@
We validate our theory in two ways. In fig.˜1(d), we show that our theory (solid lines) accurately predicts the average cosine similarity between tokens (averaged over all token pairs in a sequence) when propagating sequences from the TinyStories dataset (Eldan and Li, 2023) through a vanilla BERT model (Devlin et al., 2019) at initialisation (dots). In fig.˜1(e), we show the test loss of a 60-layer BERT model on TinyStories for two initialisations from each regime; models suffering rank or entropy collapse at initialisation indeed fail to train, as the trainability diagram predicts. We provide further applications of our theory in section˜5.
@@
  • We identify a sharp phase transition in the behaviour of self-attention in the limit of infinite sequence length, using tools from statistical physics (Derrida, 1981). Below a critical value $\beta_{c}$ of the initial scale $\beta$ of query/key weights, eq.˜9, attention matrices are approximately uniform, while for $\beta>\beta_{c}$, they are highly saturated, causing rank and entropy collapse, respectively, see result˜1.
  • Our results are exact in the limit of very large sequences. We calculate finite-size corrections that are essential for obtaining precise, practical predictions, see section˜3.2.
@@
  • We integrate our results on self-attention and normalisation layers with existing results on MLPs to derive simple algorithms like section˜4 that yield trainability diagrams like the one shown in fig.˜1c, see section˜4.
@@ 2 Setup and notation
We consider a vanilla Transformer encoder (Vaswani et al., 2017) that processes sequences $\{X_{t}\}_{t=1,\ldots,T}$ of $T$ tokens embedded in a $d$-dimensional space. We analyse signal propagation through a complete Transformer block comprising self-attention, residual connections, layer normalisation, and a feed-forward MLP.
Layer Norm. We consider layer normalisation (Ba et al., 2016) which centers each token embedding and rescales it by its standard deviation; for simplicity, we omit the affine transformation:
@@
Self-Attention. Given the query, key and value weight matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$, the attention score between tokens $X_{t}$ and $X_{t^{\prime}}$ is
@@
MLP. Each embedding $X_{t}$ is processed independently by a shared two-layer feed-forward network with hidden dimension $d$ and non-linearity $\phi$:
@@
Residual Connections. We write the residual connections for the self-attention and MLP blocks as
@@
where $\alpha>0$ are the strengths of the skip connections.
Sequence Structure. Since token embeddings are sampled i.i.d. at initialisation, standard concentration results (Vershynin, 2018, sec. 3.1) in high dimension $d$ imply that their norms concentrate, while pairwise overlaps scale as $O(d^{-1/2})$. Under self-attention, these tokens mix and their overlaps evolve as the sequence propagates in depth. Two key quantities to describe the evolution of token similarity are the overlap and cosine similarity matrices
which measure the degree of alignment between two tokens in a sequence. Our theory provides a framework to analytically track the evolution of these quantities through transformer blocks.
Notation.
@@ 3 A theory for signal propagation in Transformers
The main challenge in analysing information flow through a Transformer is handling the self-attention mechanism with its strong non-linearity and its normalisation step. In this section, we derive a theory for signal propagation by leveraging a formal similarity between self-attention and the Random energy model of Derrida (1981), see section˜3.1. Using tools and concepts from statistical physics, we give an asymptotically precise characterisation of signal propagation in self-attention layers in the forward pass in section˜3.2. We analyse the backward pass in section˜3.3. To complete the picture, we finally add the remaining elements of the Transformer architecture in section˜4, which yields section˜4 to compute the trainability diagram of Transformers.
@@ 3.1 Self-attention and the random energy model
The starting point of our analysis is an exact mapping of attention at initialisation to a model of disordered systems from statistical physics, called the Random Energy Model (Derrida, 1981). This mapping was previously identified by Lucibello and Mézard (2024) in the context of associative memories; our novelty lies in leveraging it to provide an exact description of signal propagation.
The Random Energy Model (REM). The REM is a simple model of a disordered system with $N$ spins $\mathbf{s}$ that has $\mathcal{O}(\exp(N))$ possible configurations. Each configuration has an energy $E(\mathbf{s})$ drawn at random from a Gaussian distribution with variance $\mathcal{O}(N)$ – hence the name Random Energy Model. The probability of each state is given by the Boltzmann distribution with inverse temperature $\beta>0$:
@@
Mapping of Self-Attention to the REM. We initialise the query/key projection parameters i.i.d. with variance $\sigma_{Q}^{2}=\sigma_{K}^{2}=\nicefrac{{\sigma_{a}}}{{d}}$ so that the attention scores, defined in eq.˜2, are normally distributed with mean 0 and variance $\sigma_{a}^{2}$. While both REM energies and self-attention scores $a_{tt^{\prime}}$ are Gaussian, the key difference between them is that attention scores are correlated. Averaging over the realisation of $W_{Q},W_{K}$, we get the following correlation structure (see section˜B.1):
$\mathbb{E}\left[a_{ts}a_{\tau\sigma}\right]=\sigma_{Q}^{2}\sigma_{K}^{2}\,(X_{t}\cdot X_{\tau})(X_{s}\cdot X_{\sigma})=\sigma_{a}^{2}q_{t\tau}q_{s\sigma}.$ (8)
We extend the REM framework to account for these correlations; consequently, the critical inverse temperature derived in result˜1 depends on these underlying geometric quantities.
@@
Scaling of Query/Key Weights. In MLPs, the natural scaling for weight variance is $\mathcal{O}(1/d)$ (He et al., 2015), ensuring that neuron pre-activations remain $\mathcal{O}(1)$. For self-attention, the REM analogy suggests that the natural scaling for query and key weight variance should make the attention score variance $\mathcal{O}(\log T)$, where $T$ is the sequence length, to match the energy of a REM with $N$ degrees of freedom, whose energy typically scales as $\mathcal{O}(N)$ with fluctuations of $\mathcal{O}(\sqrt{N})$, while the partition sum runs over $\mathcal{O}(e^{N})$ configurations. Self-attention only sums over $T$ tokens in the sequence, suggesting that attention scores should scale like $\mathcal{O}(\sqrt{\log T})$. This motivates us to control the variance of attention scores via a constant inverse "temperature" parameter $\beta\in\mathbb{R}$ as:
@@
This scaling is, to our knowledge, novel and complements the existing literature on various reparametrisations of the infinite-width limit (Yang and Hu, 2020) by accounting for the infinite sequence length limit.
Figure 2: Phase diagram for a single layer of self-attention. We use result˜1 to plot the average cosine similarity between pairs of tokens after one layer of self-attention as a function of the query/key variance parameter $\beta$ and the input average cosine similarity $\rho$. (Left): Theoretical phase diagram obtained from result˜1 (with $q=1$ and $p=\rho$). For $\beta<\beta_{c}$, we observe a rank collapse phase, where all input tokens map to a single output direction and the cosine similarity saturates at 1. For $\beta>\beta_{c}$, token diversity is preserved, but entropy collapse emerges. (Right): Simulations with embedding dimension $d=512$ and sequence length $T=1024$ qualitatively reproduce the theoretical transition, with deviations attributed to finite-size effects, as discussed in section˜3.2.
@@ 3.2 Forward Signal Propagation through a self-attention layer
By tracking the average squared norm $\mathbb{E}\langle q_{tt}\rangle:=q$ and the average overlap $\mathbb{E}\langle q_{ts}\rangle:=p$, we describe the evolution of the average cosine similarity $\rho$ and of the inverse participation ratio (IPR), which serves as an indicator of entropy collapse. The IPR of an attention row is defined for all queries $t\in T$ and $r\in\mathbb{N}$ as
@@
For $r=2$, the IPR estimates the effective number of tokens receiving attention: a small IPR, scaling as $o_{T}(1)$, corresponds to delocalised attention, while a non-vanishing IPR indicates that only $O_{T}(1)$ keys matter, signalling sparse self-attention and entropy collapse. Larger values of $r$ sharpen the distinction between localised and delocalised attention vectors.
We state the result below and provide its derivation in section˜B.2.
@@
Let $W_{Q}$ and $W_{K}$ be initialised with i.i.d. entries with variance $\sigma^{2}_{Q}=\sigma^{2}_{K}=\beta\sqrt{\log T}/d$, and $W_{V}$ with variance $\sigma_{V}^{2}=\sigma_{v}^{2}/d$. For a sequence with average squared token norm $q$ and average pairwise overlap $p$, define the critical initialisation scale $\beta_{c}(q,p)\equiv\sqrt{\frac{2}{q(q-p)}}$. In the limit of infinite width $d\to\infty$ and infinite sequence length $T\to\infty$, we then have that:
@@
Definition of $\beta_{c}$ and Finite-Size Effects.
A key element of our analysis is the convergence behaviour of the softmax normalisation $Z(\beta)=\sum_{\tau=1}^{T}e^{a_{t\tau}}$, which depends on $\beta$ through the scores $a_{t\tau}$ (see eq.˜2). We define $\beta_{c}$ as the value of $\beta$ at which the law of large numbers for $Z$ breaks down, i.e. the value for which $Z(\beta)$ no longer concentrates around its mean as $T\to\infty$. However, as noted by Ben Arous et al. (2005), the central limit theorem for ZZ—which guarantees Gaussian fluctuations around $\mathbb{E}Z$—already fails at a lower threshold $\tilde{\beta}_{c}=\beta_{c}/2$. Since $T$ is large but finite in practice, we therefore expect our asymptotic predictions to remain accurate for $\beta<\tilde{\beta}_{c}$, where fluctuations are still approximately Gaussian. For $\beta>\tilde{\beta}_{c}$, these finite-size fluctuations become non-Gaussian, signalling a gradual crossover toward the broken-LLN regime rather than a sharp transition exactly at $\beta_{c}$. This mechanism accounts for the discrepancies between theory and simulation observed near the critical inverse temperature in fig.˜2.
Rank Collapse. The phase diagram shown in fig.˜2 exhibits a phase transition in the cosine similarity between tokens after self-attention. For small initialisation $\beta<\beta_{c}$, self-attention operates in a ‘spread attention’ phase where the attention layer effectively outputs the average of all tokens, making the pair-wise cosine similarity between tokens saturate at one, resulting in a rank-one representation matrix. In the absence of skip connections, this behaviour leads to rank collapse, which makes the Transformer untrainable. This behaviour is also reminiscent of the “clustering” property analysed by Geshkovski et al. (2023b); Bruno et al. (2025); Chen et al. (2025).
Entropy Collapse.For $\beta>\beta_{c}$, the self-attention layer preserves diversity, and hence information, among input tokens, suggesting this regime as a viable initialisation. However, the non-vanishing value of the IPR for $\beta>\beta_{c}$, eq.˜12, reveals that self-attention is in a localised phase in this regime: it only attends to a few tokens which are determined by initialisation, rather than learnt, leading to the training instabilities observed by Zhai et al. (2023). In other words, for $\beta>\beta_{c}$ self-attention suffers from entropy collapse, which cannot be avoided using skip connections. This behaviour is a central novelty of our analysis: Noci et al. (2022) assumes the model operates entirely in the rank-collapse phase, thereby overlooking this distinct failure mode, while the “annealed” approximation of Cowsik et al. (2024) does not capture the large-deviation behaviour underlying entropy collapse.